LPC meeting summary 02-07-2018 - final

Minutes overview      LPC home

Minutes and Summary

Main purpose of the meeting: Summarize the status of the on-going special run

Introduction (Christoph Schwick):

Recovery after TS1 a bit rocky. Various issues from slides.

Started with realignment of roman pots and TCT in IP1 which all had interventions. Calibration data taken for 90m run as part of realignment cycle. Had to validate 77 lossmaps for three different cycles. Needed 15 ramps (9 ramps minimum). Originally plan was to ramp up 90m Thursday only starting today, so four day delay. Rather complicated programme and some delays, for various issues - take from slide. Mostly not
related to physics program. ALICE ++ could not be used with full. ALICE would like to stay at -- for future physics running, which is ideal for machine, but if needed machine can prepare also for ++ with a bit of heads up. CMS still have issue with HCAL endcap, but not critical for 90m run which will continue.

For the VdM programme. ATLAS was allowed by MPP to separate both planes to get luminosity down to a negligible luminosity. IP2/8 scan went smoothly. For IP1/5 scan, CMS got firealarm which left detector off for 8 hours and required redoing full scan, but managed to do so within existing fill. Had to cancel sandwich fill after two failed attempts in order to get 90m run started. ATLAS observed large horizontal orbit drifts - Jorg Wenninger: Expected as orbit feedback is off and scan happen during maximal tide forces. VdM Data for all experiments look good so far.

90m ramp-up has started. 72b bunch just completed 2 hours running, now filling 300b. Beam parameters look reasonable. Will keep 734b fill for ~10 hours and then try to switch 50ns.  Beam emittances are around
1.6mum.  For 50ns programme, if TOTEM needs to reduce pile-up, would need to separate. Ideally in separation plane, but is in vertical which physics not too happy. Could also separate in crossing plane, but only in one direction. Have asked for next fill to reduce bunch intensity to 0.7-0.8 instead of 0.9 for first fill. For now have used about 54 hours for preparation of 90m, leaving 5 days for physics programme.

Plans after 90m. ATLAS and CMS would like to have high lumi fill with 2536 or 2460 scheme depending on availability of expert to change AGK.Would be kept as long as normal (~13 hours) Do not foresee need to do intensity ramp-up after 90m run. Will not enable collisions in IP2/8 as full validation was done with beams separated in IP2/8. ATLAS will do low-mu run in agreement with CMS which will aim for 20 hours with mu=2 in ATLAS. Will also do BRST calibration soon after 90m programme, RF cogging to shift IP with 5mm. Also request for test new PPLP ramp - MD or not? In past worked for normal filling schemes - to be seen if this can be made fully transparent.

During ATLAS/CMS scan - filling scheme was not transferred correctly by DIP and had to be reapplied. LDM data was also not available at start of ATLAS/CMS fill, but was fixed quickly from remote. Witold suggest that ALICE/LHCb checks this for their fill.